artlover03:

Use the quadratic formula to solve for x. 2x^2+5x=4 Round your answer to the nearest hundredth. If there is more than one solution, separate them with commas.

jhonyy9:

@supie any idea ?

artlover03:

i got to\[-5\pm \sqrt{57} \div4\] WHY DID THAT TAKE ME SO LONG TO WRITE

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 i got to\[-5\pm \sqrt{57} \div4\] WHY DID THAT TAKE ME SO LONG TO WRITE \(\color{#0cbb34}{\text{End of Quote}}\) but this you dont written it right for numerator you need using parentheses

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 i got to\[(-5\pm \sqrt{57}) \div4\] WHY DID THAT TAKE ME SO LONG TO WRITE \(\color{#0cbb34}{\text{End of Quote}}\) but this you dont written it right for numerator you need using parentheses \(\color{#0cbb34}{\text{End of Quote}}\)

artlover03:

wydm? i cant even solve the rest i need help lol

AZ:

\( x = \dfrac{ -b \pm \sqrt{b^2 - 4ac}}{2a} \) when the equation is \( ax^2 + bx + c = 0\) so your equation is \( 2x^ 2 + 5x - 4 = 0\) and so yeah, you do get \( x = \dfrac{ -5 \pm \sqrt{57}}{4} \)

AZ:

So you're correct so far Let's separate the answers. Do you have a calculator with you?

AZ:

\(\pm\) means plus or minus so we have \( x = \dfrac{ -5 + \sqrt{57}}{4} \) and \( x = \dfrac{ -5 - \sqrt{57}}{4} \) you'll need to use a calculator to find the values of it

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \( x = \dfrac{ -b \pm \sqrt{b^2 - 4ac}}{2a} \) when the equation is \( ax^2 + bx + c = 0\) so your equation is \( 2x^ 2 + 5x - 4 = 0\) and so yeah, you do get \( x_1,2 = \dfrac{ -5 \pm \sqrt{57}}{4} \) \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \( x = \dfrac{ -b \pm \sqrt{b^2 - 4ac}}{2a} \) when the equation is \( ax^2 + bx + c = 0\) so your equation is \( 2x^ 2 + 5x - 4 = 0\) and so yeah, you do get \( x_1,2 = \dfrac{ -5 \pm \sqrt{57}}{4} \) \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\) in this way written is right with x_1,2 bc. there are two roots

artlover03:

+= .6374... -= -3.1374...

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 in this way written is right with x_1,2 bc. there are two roots \(\color{#0cbb34}{\text{End of Quote}}\) Yes, agreed

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 in this way written is right with x_1,2 bc. there are two roots \(\color{#0cbb34}{\text{End of Quote}}\) Yes, agreed \(\color{#0cbb34}{\text{End of Quote}}\) explained perfect - good job

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @artlover03 += .6374... -= -3.1374... \(\color{#0cbb34}{\text{End of Quote}}\) Good job! So those are your two values for x Make sure to round to the nearest hundredth

artlover03:

which would be += 0.64 -= -3.14 ?

AZ:

Correct!

artlover03:

OMG! thank you so much, have a good one!

AZ:

You too!